OpenStudy (anonymous):

change to logarithmic form y=3^x

OpenStudy (anonymous):

would that be \[\log _{3}x=y\]

hartnn (hartnn):

that would give you x = 3^y

hartnn (hartnn):

see, \(\large a^b =c \implies b = \log_ac\)

hartnn (hartnn):

so, try again ?

OpenStudy (anonymous):

would it be \[x=\log_3y\]

hartnn (hartnn):

absolutely correct!

OpenStudy (anonymous):

thank you for your help :)

hartnn (hartnn):

you're welcome ^_^